OpenStudy (anonymous):

b2-2b=24 solve by completing the square

jimthompson5910 (jim_thompson5910):

start off by taking half of -2 (the b term), then squaring it (-2/2)^2 = (-1)^2 = 1

jimthompson5910 (jim_thompson5910):

you then add 1 to both sides b^2 - 2b + 1 = 24+1 b^2 - 2b + 1 = 25 the left side is now a perfect square, so it factors to (b-1)^2 giving you (b-1)^2 = 25

jimthompson5910 (jim_thompson5910):

I'm sure you see how to finish up